# Information Security Based on LLM Approaches: A Review

### v1

@@ 1. Introduction
With the rapid development of information technology, the issue of information security has become increasingly prominent, posing a common challenge for individuals, enterprises, and even countries. Threats such as leakage of personal information, loss of financial data, and network fraud are increasing; enterprises are facing hidden dangers such as database security, system vulnerability, and intellectual property protection; and at the national level, cyberattacks and information wars are seriously jeopardizing the security of critical infrastructure and information systems (shen2007survey, ). Therefore, improving the security protection capability of information systems has become a crucial issue that needs to be addressed urgently. In recent years, as artificial intelligence, particularly Large Language Models (LLMs), has made significant progress in the field of natural language processing, representative models such as GPT and BERT have been widely utilized in tasks including text generation, translation, and analysis. Its powerful language understanding and generation capabilities have also shown broad application prospects in the field of information security, such as malicious code recognition, network attack detection, and vulnerability analysis(li2024stateguard, ). At the same time, the safe and effective application of large language models in the field of information security has become a new direction of study (siponen2007review, ).
The purpose of this paper is to systematically explore the current status of the application of large language models in information security, the challenges it faces, and its practical effects. Through the combination of literature review, case study, and empirical research, we assess the advantages and shortcomings of the Large Language Model in enhancing information security capabilities and provide theoretical support and practical guidance for its further development in the field of information security.
@@ 2.1. Neural Network
In the field of artificial intelligence, a neural network is a mathematical model that simulates the information transfer mechanism of human neurons. Its basic structure includes an input layer, a hidden layer, and an output layer. The input layer is used to receive the original data and pass it to the subsequent network; the output layer generates the final prediction results; the hidden layer is located between the two, responsible for nonlinear mapping and feature extraction of the data, and the number of layers and nodes can be flexibly adjusted according to the task requirements; generally speaking, the deeper the hidden layer is, the stronger the model’s expressive ability is. Meanwhile, the training process of neural networks mainly includes two phases: forward propagation and back propagation(li2024defitail, ). In the forward propagation process, the input data passes through the neurons of each layer in turn, and is processed by the weighted summation and activation function to generate the final output; in the back propagation stage, the error between the prediction result and the real label is calculated, and the algorithms such as gradient descent are used to adjust the network weights in the reverse directparticularlytimize the model performance continuously. Costing multiple iterations of training, the model error is generally reduced until convergence(liu2017survey, ). The structure of the neural network is shown in Fig. 1
@@ 2.2. Machine Learning
  • Supervised Learning: Supervised learning relies on labeled datasets for training, i.e., each sample contains known inputs with corresponding output targets. The model gradually optimizes the parameters and improves the prediction accuracy by continuously comparing the prediction results with the real labels. The method is widely used in scenarios such as spam recognition and credit scoring. However, the limitation of supervised learning is the high cost of acquiring labeled data, which relies on a large amount of manual labeling.
@@
  • Unsupervised Learning: Unsupervised learning deals with unlabeled data to discover underlying structures or patterns in the data, such as clustering and dimensionality reduction. The model automatically recognizes the intrinsic distributional features of the data without external guidance, and is commonly used for tasks such as customer segmentation and anomaly detection(liu2025sok, ).
@@
  • Semi-supervised Learning: Semi-supervised learning combines a small amount of labeled data with a large amount of unlabeled data for training, taking into account the advantages of supervised and unsupervised methods(li2024guardians, ). In practical applications, it is particularly suitable for scenarios with high labeling costs but a large amount of data, which effectively improves the model’s generalization ability and efficiency.
@@
  • Reinforcement Learning: Reinforcement learning is centered on the interaction of an intelligent body (agent) with the environment to maximize long-term rewards through trial-and-error mechanisms. The agent performs actions based on its current state, receives feedback (in the form of reward or punishment) from the environment, and adjusts its strategy accordingly to improve overall performance. This approach is widely used in the fields of automatic control, game strategy, and robot navigation, and is one of the important paths to realize general artificial intelligence. As shown in Fig. 2, the basic framework of reinforcement learning includes the key elements of state, action, strategy, and reward(zhou2018brief, ).
@@
  • Deep Learning: Deep learning is a machine learning method based on multi-layer neural networks, capable of hierarchical feature extraction and abstract representation of data. It has achieved remarkable results in the fields of image recognition, speech processing, and natural language understanding. The construction of large language models is highly dependent on deep learning, and by training on large-scale corpora, the models can capture complex linguistic structure and semantic information, and realize tasks such as syntactic analysis, context understanding, and text generation. In addition, deep learning also provides fundamental support for the application of large language models in information security. By training neural networks to identify potential threat patterns, the model can be used for security tasks such as malicious behavior detection, abnormal traffic identification, and data leakage prevention and control, showing significant potential in enhancing cyberspace governance capabilities(pouyanfar2018survey, ).
@@ 2.3. Large Model
With the continuous advancement of deep learning and computational resources, large model technology has achieved breakthroughs in the fields of natural language processing, computer vision, and speech recognition. Its development began with early artificial neural networks and back propagation algorithms, and the deep belief network proposed by Hinton in 2006 promoted the rise of deep learning, while the success of AlexNet in 2012 triggered the widespread application of deep learning. Subsequently, visual models such as VGG and ResNet have continuously optimized the network structure and improved the model performance(chen2018system, ). In the field of natural language processing, pre-trained language models based on the Transformer architecture, such as BERT and GPT, have been introduced one after another, promoting a leap in language comprehension and generation capabilities. Meanwhile, systems such as AlphaGo, which incorporate multiple technologies, have also verified the potential of large models in complex intelligence tasks. Currently, the development of large models is supported by key technologies such as pre-training and fine-tuning mechanisms, distributed training, model compression, and Generative Adversarial Networks (GANs), which have become an important engine for the development of AI(zhao2023survey, ).
@@
Large-scale pretrained language models (PLMs) are natural language understanding and generation models trained by deep learning methods using large-scale unlabeled corpora. These models usually adopt a “pre-training + fine-tuning” strategy: the pre-training phase learns linguistic features on general-purpose text, while the fine-tuning phase optimizes on labeled data according to specific tasks. The common structure is based on the Transformer framework, which consists of two main parts: the Encoder and the Decoder, which are used to model the contextual relationship between input and output sequences(zou2025malicious, ). The current mainstream models can be roughly categorized into two types: autoregressive models (e.g., GPT), which predict the next word through the sequence antecedent, and self-encoding models (e.g., BERT), which use a masking mechanism to capture bidirectional contextual semantics. The former emphasizes text generation capabilities, while the latter is more suitable for comprehension-type tasks. Model training often combines a cross-entropy loss function with multi-task learning (e.g., MLM, NSP), designed to enhance semantic expressiveness and migration generalization. Pre-trained language models have been widely used in Q&A systems, text categorization, code generation, and information security scenarios, and have become the core supporting technology in the current field of natural language processing(bharathi2024analysis, ).
@@
Language modeling based on autoregressive models: Autoregressive language models use the previous known words in the sequence as conditions to predict the next word step by step, and the whole generation process satisfies the Markov property. Typical representatives include RNNLM (Recurrent Neural Network Language Model) and Transformer-based GPT series(niu2024unveiling, ). In these models, the hidden state of each time step is determined by the information of the previous word, and the output is calculated by a Softmax layer to predict the probability. In the case of RNNLM, the training objective is to maximize the log-likelihood of each word in the entire corpus, which is expressed mathematically as follows:
@@
Included among these, $P(w_{t}\mid w_{1},\ldots,w_{t-1})$ denotes the probability of the current word under the given historical conditions. In contrast, GPT models adopt a multi-layer Transformer decoder structure and utilize a self-attention mechanism to effectively capture long-distance dependencies, thus achieving stronger context modeling capabilities. Such models perform well in tasks such as text generation and code auto-completion.
@@
where $T$ denotes the length of the input sequence, $V$ denotes the size of the vocabulary list, $y_{t,i}$ is the actual label (one-hot encoding) of the word $i$ in the $t$ time step, and $\hat{y}_{t,i}$ is the corresponding probability predicted by the model. The method can effectively capture the overall structural information of the input sequence and performs well in a variety of natural language processing tasks, especially for semantic understanding and context modeling(kasneci2023chatgpt, ).
@@ 2.4. Transformer
To enhance the capability of natural language understanding and deep semantic modeling, the Transformer architecture has been proposed. Originally designed for machine translation tasks, the basic structure of the Transformer is shown in Fig. 3. The overall structure can be divided into left and right parts: the Encoder on the left side and the Decoder on the right side. In the Encoder section, each word in the input sequence is first combined with Positional Encoding through Input Embedding to generate vector representations. These representations are then sequentially processed through multiple layers consisting of Multi-Head Attention and Feed-Forward Neural Network(zhong2023tackling, ), each equipped with Residual Connection and Layer Normalization ( Residual Connection and Layer Normalization) between each layer. The decoder structure is similar to the encoder, but introduces Masked Multi-Head Attention after Multi-Head Attention to avoid future information leakage and receives the output from the encoder for inter-attention computation. Finally, after Linear Layer and Softmax (normalized exponential function) operations, the model outputs the corresponding translation or prediction results(khan2022transformers, ).
@@ 2.5. Network Security Attacks
Network security attack refers to the attacker’s use of technical means of computer systems, network resources, or data and information for illegal access, theft, destruction, or tampering behavior(hoque2014network, ). According to the different attack methods and targets, common types of network attacks mainly include:
@@
  • Worm attack: Self-replicating malicious code that does not depend on the host program, spreads mainly through the network, can infect quickly, and puts great pressure on network bandwidth and system resources(bu2025smartbugbert, ).
@@
In the face of evolving network attack methods, there is an urgent need to take systematic security measures(ehlert2010survey, ), such as strengthening the authentication mechanism, regularly updating system patches, deploying Intrusion Detection Systems (IDS), and improving user security awareness. Meanwhile, attack detection technology also plays a key role in the network security defense system, and its core methods include:
@@
  • Behavioral Analysis: Modeling and anomaly detection based on user or system operational behavioral characteristics helps to discover hidden attacks such as Advanced Persistent Threats (APTs)(li2021hybrid, ).
@@ 2.6. Security Breach
Although common security tools such as antivirus software have a certain vulnerability repair function, with the continuous evolution of attack technology, new types of vulnerabilities continue to emerge, and their hidden complexity is also increasing(li2017discovering, ). Some vulnerabilities are hidden in the underlying structure of the system or triggered by the interaction of multiple components, which are difficult to discover and repair promptly through traditional methods. At the same time, the continuous compression of the software development cycle has also accelerated the pace from the emergence of vulnerabilities to be exploited, so that the existing detection methods face the double challenge of efficiency and accuracy.
In addition, vulnerability detection often relies on high-level technical experts and complex detection tools, which puts high demands on resource allocation and technical capabilities. Therefore, building an efficient, intelligent, and automated vulnerability detection mechanism has become a key direction to improve network defense capability. In recent years, Large Language Models (LLMs) have shown strong potential in code semantic understanding and vulnerability prediction, promoting the transformation of vulnerability detection technology to deep learning-driven intelligence, and providing a brand new way of thinking to improve system security and robustness(al2020dealing, ).
@@ 2.7. Malicious Code
Malicious code refers to malicious programs designed to disrupt system functionality or steal user data. Common types include viruses, worms, Trojan horses, and spyware. Viruses and worms can replicate themselves, with worms spreading autonomously over the network; Trojans use camouflage to lure users to execute them; and spyware focuses on surveillance behavior and information theft. According to the propagation path, malicious code can be spread through email, web pages, removable devices, and other media; according to the target of attack, the target can cover individual users, enterprise networks, and even government systems. Classification and characterization of malicious code are the basis for building detection and defense mechanisms. Behavioral features such as file operations, registry modifications, system calls, and network communications, as well as static attributes such as string patterns and structural features, are the important basis for malicious code identification(alazab2015profiling, ).
Current mainstream detection methods include static analysis and dynamic analysis. The former analyzes the structure and logic of the source code or binary file to determine the potential risk without running the program, which is suitable for preventive detection but susceptible to obfuscation and encryption techniques; the latter monitors the program behavior during runtime, which is more reflective of the real threat but may be interfered with by the evasion mechanism of malicious code on the detection environment.
@@ 2.8. Cryptography
Cryptography is the core discipline of information security, which mainly contains two aspects of cryptographic coding and cryptanalysis. Cryptographic coding typically committed to involves generating programs to protect information, encrypting plaintext with encryption algorithms decrypting ciphertext with decryption algorithms(bu2025enhancing, ). A complete cryptographic system usually contains plaintext, ciphertext, key, encryption algorithm, and decryption algorithm, and other basic elements; the three work together to ensure the confidentiality, integrity, and availability of information.
Modern cryptosystems are divided into two categories: symmetric encryption and asymmetric encryption. Symmetric encryption uses the same key to complete the encryption and decryption process, which has the advantages of fast encryption and decryption speed and high computational efficiency, and is suitable for the protection of large amounts of data. Typical symmetric encryption algorithms include AES, DES, and so on. Asymmetric encryption is based on a pair of keys - public key (public key) and private key (private key), public key for encryption, private key for decryption. The security depends on the computational complexity of the mathematical puzzle. Common asymmetric encryption algorithms include RSA, DSA, and Elliptic Curve Cryptography (ECC)(hellman2002overview, ).
Cryptographic algorithms should be designed to satisfy the correctness requirement, i.e., decrypting the encrypted ciphertext should restore the original plaintext. In addition, according to Kirchhoff’s principle, the security of a cryptosystem should depend only on the secrecy of the key; even if an attacker has all the details of the encryption algorithm, the system remains reliable as long as the key is secure. Modern cryptographic communication typically involves generating keys through key generation algorithms, encrypting plaintext with encryption algorithms, and decrypting ciphertext with decryption algorithms, thereby constituting a complete secure transmission process. The process is shown in Fig. 4.

### v2

@@ 1. Introduction
With the rapid development of information technology, the issue of information security has become increasingly prominent, posing a common challenge for individuals, enterprises, and even countries. Threats such as leakage of personal information, loss of financial data, and network fraud are increasing; enterprises are facing hidden dangers such as database security, system vulnerability, and intellectual property protection; and at the national level, cyberattacks and information wars are seriously jeopardizing the security of critical infrastructure and information systems (Shen et al., 2007). Therefore, improving the security protection capability of information systems has become a crucial issue that needs to be addressed urgently(Chang et al., 2024). In recent years, as artificial intelligence, particularly Large Language Models (LLMs), has made significant progress in the field of natural language processing, representative models such as GPT and BERT have been widely utilized in tasks including text generation, translation, and analysis. Its powerful language understanding and generation capabilities have also shown broad application prospects in the field of information security, such as malicious code recognition, network attack detection, and vulnerability analysis(Li et al., 2024f). At the same time, the safe and effective application of large language models in the field of information security has become a new direction of study (Siponen and Oinas-Kukkonen, 2007).
The purpose of this paper is to systematically explore the current status of the application of large language models in information security, the challenges it faces, and its practical effects(Al-Balushi et al., 2024). Through the combination of literature review, case study, and empirical research, we assess the advantages and shortcomings of the Large Language Model in enhancing information security capabilities and provide theoretical support and practical guidance for its further development in the field of information security(Naveed et al., 2025).
@@ 2.1. Neural Network
In the field of artificial intelligence, a neural network is a mathematical model that simulates the information transfer mechanism of human neurons. Its basic structure includes an input layer, a hidden layer, and an output layer. The input layer is used to receive the original data and pass it to the subsequent network; the output layer generates the final prediction results(Zhao et al., 2024); the hidden layer is located between the two, responsible for nonlinear mapping and feature extraction of the data, and the number of layers and nodes can be flexibly adjusted according to the task requirements(Zhao and Zhang, 2024); generally speaking, the deeper the hidden layer is, the stronger the model’s expressive ability is. Meanwhile, the training process of neural networks mainly includes two phases: forward propagation and back propagation(Li et al., 2024c). In the forward propagation process, the input data passes through the neurons of each layer in turn, and is processed by the weighted summation and activation function to generate the final output; in the back propagation stage, the error between the prediction result and the real label is calculated, and the algorithms such as gradient descent are used to adjust the network weights in the reverse directparticularlytimize the model performance continuously. Costing multiple iterations of training, the model error is generally reduced until convergence(Liu et al., 2017). The structure of the neural network is shown in Fig. 1
@@ 2.2. Machine Learning
  • Supervised Learning: Supervised learning relies on labeled datasets for training, i.e., each sample contains known inputs with corresponding output targets(Igual and Seguí, 2024). The model gradually optimizes the parameters and improves the prediction accuracy by continuously comparing the prediction results with the real labels. The method is widely used in scenarios such as spam recognition and credit scoring. However, the limitation of supervised learning is the high cost of acquiring labeled data, which relies on a large amount of manual labeling(Lim et al., 2024).
@@
  • Unsupervised Learning: Unsupervised learning deals with unlabeled data to discover underlying structures or patterns in the data, such as clustering and dimensionality reduction. The model automatically recognizes the intrinsic distributional features of the data without external guidance, and is commonly used for tasks such as customer segmentation and anomaly detection(Liu and Li, 2025).
@@
  • Semi-supervised Learning: Semi-supervised learning combines a small amount of labeled data with a large amount of unlabeled data for training, taking into account the advantages of supervised and unsupervised methods(Li et al., 2024e). In practical applications, it is particularly suitable for scenarios with high labeling costs but a large amount of data, which effectively improves the model’s generalization ability and efficiency.
@@
  • Reinforcement Learning: Reinforcement learning is centered on the interaction of an intelligent body (agent) with the environment to maximize long-term rewards through trial-and-error mechanisms. The agent performs actions based on its current state, receives feedback (in the form of reward or punishment) from the environment, and adjusts its strategy accordingly to improve overall performance(Gu et al., 2024). This approach is widely used in the fields of automatic control, game strategy, and robot navigation, and is one of the important paths to realize general artificial intelligence. As shown in Fig. 2, the basic framework of reinforcement learning includes the key elements of state, action, strategy, and reward(Zhou, 2018).
@@
  • Deep Learning: Deep learning is a machine learning method based on multi-layer neural networks, capable of hierarchical feature extraction and abstract representation of data. It has achieved remarkable results in the fields of image recognition, speech processing, and natural language understanding(Herrmann and Kollmannsberger, 2024). The construction of large language models is highly dependent on deep learning, and by training on large-scale corpora, the models can capture complex linguistic structure and semantic information, and realize tasks such as syntactic analysis, context understanding, and text generation(Chen and Chen, 2026). In addition, deep learning also provides fundamental support for the application of large language models in information security. By training neural networks to identify potential threat patterns, the model can be used for security tasks such as malicious behavior detection, abnormal traffic identification, and data leakage prevention and control, showing significant potential in enhancing cyberspace governance capabilities(Pouyanfar et al., 2018).
@@ 2.3. Large Model
With the continuous advancement of deep learning and computational resources, large model technology has achieved breakthroughs in the fields of natural language processing, computer vision, and speech recognition(Zhu et al., 2024). Its development began with early artificial neural networks and back propagation algorithms, and the deep belief network proposed by Hinton in 2006 promoted the rise of deep learning, while the success of AlexNet in 2012 triggered the widespread application of deep learning. Subsequently, visual models such as VGG and ResNet have continuously optimized the network structure and improved the model performance(Chen et al., 2018). In the field of natural language processing, pre-trained language models based on the Transformer architecture, such as BERT and GPT, have been introduced one after another, promoting a leap in language comprehension and generation capabilities. Meanwhile, systems such as AlphaGo, which incorporate multiple technologies, have also verified the potential of large models in complex intelligence tasks. Currently, the development of large models is supported by key technologies such as pre-training and fine-tuning mechanisms, distributed training, model compression, and Generative Adversarial Networks (GANs), which have become an important engine for the development of AI(Zhao et al., 2023).
@@
Large-scale pretrained language models (PLMs) are natural language understanding and generation models trained by deep learning methods using large-scale unlabeled corpora. These models usually adopt a “pre-training + fine-tuning” strategy: the pre-training phase learns linguistic features on general-purpose text, while the fine-tuning phase optimizes on labeled data according to specific tasks. The common structure is based on the Transformer framework, which consists of two main parts: the Encoder and the Decoder, which are used to model the contextual relationship between input and output sequences(Zou et al., 2025). The current mainstream models can be roughly categorized into two types: autoregressive models (e.g., GPT), which predict the next word through the sequence antecedent, and self-encoding models (e.g., BERT), which use a masking mechanism to capture bidirectional contextual semantics. The former emphasizes text generation capabilities, while the latter is more suitable for comprehension-type tasks. Model training often combines a cross-entropy loss function with multi-task learning (e.g., MLM, NSP), designed to enhance semantic expressiveness and migration generalization. Pre-trained language models have been widely used in Q&A systems, text categorization, code generation, and information security scenarios, and have become the core supporting technology in the current field of natural language processing(Bharathi Mohan et al., 2024).
@@
Language modeling based on autoregressive models: Autoregressive language models use the previous known words in the sequence as conditions to predict the next word step by step, and the whole generation process satisfies the Markov property. Typical representatives include RNNLM (Recurrent Neural Network Language Model) and Transformer-based GPT series(Niu et al., 2024). In these models, the hidden state of each time step is determined by the information of the previous word, and the output is calculated by a Softmax layer to predict the probability. In the case of RNNLM, the training objective is to maximize the log-likelihood of each word in the entire corpus, which is expressed mathematically as follows:
@@
Included among these, $P(w_{t}\mid w_{1},\ldots,w_{t-1})$ denotes the probability of the current word under the given historical conditions. In contrast, GPT models adopt a multi-layer Transformer decoder structure and utilize a self-attention mechanism to effectively capture long-distance dependencies, thus achieving stronger context modeling capabilities. Such models perform well in tasks such as text generation and code auto-completion(Zhou et al., 2024).
@@
where $T$ denotes the length of the input sequence, $V$ denotes the size of the vocabulary list, $y_{t,i}$ is the actual label (one-hot encoding) of the word $i$ in the $t$ time step, and $\hat{y}_{t,i}$ is the corresponding probability predicted by the model. The method can effectively capture the overall structural information of the input sequence and performs well in a variety of natural language processing tasks, especially for semantic understanding and context modeling(Kasneci et al., 2023).
@@ 2.4. Transformer
To enhance the capability of natural language understanding and deep semantic modeling, the Transformer architecture has been proposed. Originally designed for machine translation tasks, the basic structure of the Transformer is shown in Fig. 3. The overall structure can be divided into left and right parts: the Encoder on the left side and the Decoder on the right side. In the Encoder section, each word in the input sequence is first combined with Positional Encoding through Input Embedding to generate vector representations. These representations are then sequentially processed through multiple layers consisting of Multi-Head Attention and Feed-Forward Neural Network(Zhong et al., 2023), each equipped with Residual Connection and Layer Normalization ( Residual Connection and Layer Normalization) between each layer. The decoder structure is similar to the encoder, but introduces Masked Multi-Head Attention after Multi-Head Attention to avoid future information leakage and receives the output from the encoder for inter-attention computation. Finally, after Linear Layer and Softmax (normalized exponential function) operations, the model outputs the corresponding translation or prediction results(Khan et al., 2022).
@@ 2.5. Network Security Attacks
Network security attack refers to the attacker’s use of technical means of computer systems, network resources, or data and information for illegal access, theft, destruction, or tampering behavior(Hoque et al., 2014). According to the different attack methods and targets, common types of network attacks mainly include:
@@
  • Worm attack: Self-replicating malicious code that does not depend on the host program, spreads mainly through the network, can infect quickly, and puts great pressure on network bandwidth and system resources(Bu et al., 2025b).
@@
In the face of evolving network attack methods, there is an urgent need to take systematic security measures(Ehlert et al., 2010), such as strengthening the authentication mechanism, regularly updating system patches, deploying Intrusion Detection Systems (IDS), and improving user security awareness. Meanwhile, attack detection technology also plays a key role in the network security defense system, and its core methods include:
@@
  • Behavioral Analysis: Modeling and anomaly detection based on user or system operational behavioral characteristics helps to discover hidden attacks such as Advanced Persistent Threats (APTs)(Li and others, 2021).
@@ 2.6. Security Breach
Although common security tools such as antivirus software have a certain vulnerability repair function, with the continuous evolution of attack technology, new types of vulnerabilities continue to emerge, and their hidden complexity is also increasing(Li et al., 2017). Some vulnerabilities are hidden in the underlying structure of the system or triggered by the interaction of multiple components, which are difficult to discover and repair promptly through traditional methods. At the same time, the continuous compression of the software development cycle has also accelerated the pace from the emergence of vulnerabilities to be exploited, so that the existing detection methods face the double challenge of efficiency and accuracy(Li et al., 2024a).
In addition, vulnerability detection often relies on high-level technical experts and complex detection tools, which puts high demands on resource allocation and technical capabilities. Therefore, building an efficient, intelligent, and automated vulnerability detection mechanism has become a key direction to improve network defense capability. In recent years, Large Language Models (LLMs) have shown strong potential in code semantic understanding and vulnerability prediction, promoting the transformation of vulnerability detection technology to deep learning-driven intelligence, and providing a brand new way of thinking to improve system security and robustness(Al Naeem et al., 2020).
@@ 2.7. Malicious Code
Malicious code refers to malicious programs designed to disrupt system functionality or steal user data. Common types include viruses, worms, Trojan horses, and spyware. Viruses and worms can replicate themselves, with worms spreading autonomously over the network; Trojans use camouflage to lure users to execute them; and spyware focuses on surveillance behavior and information theft. According to the propagation path, malicious code can be spread through email, web pages, removable devices, and other media; according to the target of attack, the target can cover individual users, enterprise networks, and even government systems. Classification and characterization of malicious code are the basis for building detection and defense mechanisms. Behavioral features such as file operations, registry modifications, system calls, and network communications, as well as static attributes such as string patterns and structural features, are the important basis for malicious code identification(Alazab, 2015).
Current mainstream detection methods include static analysis and dynamic analysis. The former analyzes the structure and logic of the source code or binary file to determine the potential risk without running the program, which is suitable for preventive detection but susceptible to obfuscation and encryption techniques; the latter monitors the program behavior during runtime, which is more reflective of the real threat but may be interfered with by the evasion mechanism of malicious code on the detection environment(Patsakis et al., 2024).
@@ 2.8. Cryptography
Cryptography is the core discipline of information security, which mainly contains two aspects of cryptographic coding and cryptanalysis. Cryptographic coding typically committed to involves generating programs to protect information, encrypting plaintext with encryption algorithms decrypting ciphertext with decryption algorithms(Bu et al., 2025a). A complete cryptographic system usually contains plaintext, ciphertext, key, encryption algorithm, and decryption algorithm, and other basic elements; the three work together to ensure the confidentiality, integrity, and availability of information(Subramani et al., 2025).
Modern cryptosystems are divided into two categories: symmetric encryption and asymmetric encryption. Symmetric encryption uses the same key to complete the encryption and decryption process, which has the advantages of fast encryption and decryption speed and high computational efficiency, and is suitable for the protection of large amounts of data. Typical symmetric encryption algorithms include AES, DES, and so on. Asymmetric encryption is based on a pair of keys - public key (public key) and private key (private key), public key for encryption, private key for decryption. The security depends on the computational complexity of the mathematical puzzle. Common asymmetric encryption algorithms include RSA, DSA, and Elliptic Curve Cryptography (ECC)(Hellman, 2002).
Cryptographic algorithms should be designed to satisfy the correctness requirement, i.e., decrypting the encrypted ciphertext should restore the original plaintext. In addition, according to Kirchhoff’s principle, the security of a cryptosystem should depend only on the secrecy of the key; even if an attacker has all the details of the encryption algorithm, the system remains reliable as long as the key is secure(Radanliev, 2024). Modern cryptographic communication typically involves generating keys through key generation algorithms, encrypting plaintext with encryption algorithms, and decrypting ciphertext with decryption algorithms, thereby constituting a complete secure transmission process. The process is shown in Fig. 4.